# Rethinking the Evaluation of Out-of-Distribution Detection: A Sorites Paradox

### v1

@@ 1 Introduction
Figure 1: Examples of images from IS-OOD benchmark. ImageNet-21K is divided into subsets with different semantic and covariate shift levels relative to ImageNet-1K. As semantic shift increases, images of the subsets change from marginal samples (such as animal subspecies) to more distinct OOD categories (such as "gasket"). As covariate shift increases, the covariate contents like styles transition from object-centered photos to advertisements or monochrome images.
@@ 2.4 Metrics
$correlation=\frac{\sum_{1}^{n}(x_{i}-\bar{x})(i-\frac{n+1}{2})}{\sqrt{\sum_{1}^{n}(x_{i}-\bar{x})^{2}\sum_{1}^{n}(i-\frac{n+1}{2})^{2}}},$ (5)
@@
$sensitivity=\left|\frac{\sum_{1}^{n}(x_{i}-\bar{x})(i-\frac{n+1}{2})}{\sum_{1}^{n}(i-\frac{n+1}{2})^{2}}\right|.$ (6)
@@ 3 Experiments
In this section, we present the experimental results and corresponding findings from the proposed benchmark. During the experiments, each time we choose one divided subset from IS-OOD as the OOD data and mix it with ImageNet-1K test set to evaluate the OOD detection models trained on ImageNet-1K training set. Since the training data are from ImageNet-1K, all the shift levels mentioned in the experiments are the shifts of the test samples relative to ImageNet-1K. The details of the evaluated OOD detection methods are introduced in Appendix E. Due to space limitations, only part of the results based on the AUROC metric are presented in this section. Complete experimental results including the metrics FPR@95 and AUPR can be found in Appendix F.
@@ 3.2 Results on Syn-IS
On Syn-IS dataset, the performance of GradNorm declines significantly. We believe this may be due to the datasets generated by text-to-image models having more prominent high-level information like the semantic contents, and methods like GradNorm may struggle to extract effective information for OOD detection from these high-level semantic contents, leading to a decline in performance. On the contrary, other OOD detection methods like MSP can benefit from this high-level information and obtain performance improvements.
@@ 4 Conclusion and Discussion
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix.
@@ Appendix C Details of the Subsets Division
As observed in Figure 9, subsets with high semantic shifts and low covariate shifts (or vice versa) contain few images, sometimes even none. This indicates that there is a certain correlation between semantic and covariate information in ImageNet-21K. Therefore, we have to sacrifice the results on some subsets (marking the results of subsets with too little data as "N/A"), in which way the levels of shift can cover a broader range of shift degrees (a greater shift degree in the highest shift level).
@@ F.5 AUPR Results on Syn-IS
Appendix G Computational Resources
@@
Appendix H Definition of Semantic and Covariate Shift
It’s worth noting that we observe ImageNet-21K labels include labels such as beach and snow, which might serve as backgrounds (covariate contents) in other datasets, but appear as semantic contents in ImageNet-21K.
We treat all the labels in ImageNet-21K (such as backgrounds) as semantic contents in the proposed IS-OOD, which might differ from some other benchmarks. In our experiments, the covariate contents only include styles, image augmentations, and lighting changes. However, future researchers can modify the collection of covariate texts according to their needs to study different elements as the covariate components.
Appendix I Data Sheet
@@ I.3 Collection Process
I.4 Preprocessing/cleaning/labeling

### v3

@@ 1 Introduction
Figure 1: Examples of images from IS-OOD benchmark. ImageNet-21K is divided into subsets with different semantic and covariate shift levels relative to ImageNet-1K. As semantic shift increases, images of the subsets change from marginal samples (such as animal subspecies) to more distinct OOD categories (such as "gasket"). As covariate shift increases, the covariate contents transition from object-centered real photos to synthetic images, and from high-definition color images to low-resolution monochrome images.
@@ 2.4 Metrics
$correlation=\frac{\sum_{i=1}^{n}(x_{i}-\bar{x})(i-\frac{n+1}{2})}{\sqrt{\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}\sum_{i=1}^{n}(i-\frac{n+1}{2})^{2}}},$ (5)
@@
$sensitivity=\left|\frac{\sum_{i=1}^{n}(x_{i}-\bar{x})(i-\frac{n+1}{2})}{\sum_{i=1}^{n}(i-\frac{n+1}{2})^{2}}\right|.$ (6)
@@ 3 Experiments
In this section, we present the experimental results and corresponding findings from the proposed benchmark. During the experiments, each time we choose one divided subset from IS-OOD as the OOD data and mix it with ImageNet-1K test set to evaluate the OOD detection models trained on ImageNet-1K training set. Since the training data are from ImageNet-1K, all the shift levels mentioned in the experiments are the shifts of the test samples relative to ImageNet-1K. All evaluated OOD detection methods are implemented using a ResNet-50 [29] classifier trained on ImageNet-1K. The details of the evaluated OOD detection methods are introduced in Appendix E. Due to space limitations, only part of the results based on the AUROC metric are presented in this section. Complete experimental results including the metrics FPR@95 and AUPR can be found in Appendix F.
@@ 3.2 Results on Syn-IS
On Syn-IS dataset, the performance of GradNorm declines significantly. We further analyze the principle of GradNorm and the characteristics of the Syn-IS dataset. We find that GradNorm tends to classify samples with uniform softmax outputs as OOD samples. Since each Syn-IS image is generated with a text that only includes one given label, the image is likely to contain a single object, unlike many images in ImageNet-21K that contain multiple objects. This difference could lead to less uniform softmax outputs on Syn-IS compared to Imagenet-21K, resulting in lower OOD scores produced by GradNorm. This insight might highlight a potential limitation of the GradNorm approach.
@@ 4 Conclusion and Discussion
Acknowledgment
This work is partially supported by Natural Science Foundation of China (No. U2336213), Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0680202), Beijing Nova Program (20230484368), Suzhou Frontier Technology Research Project (No. SYG202325), and Youth Innovation Promotion Association CAS.
@@ Acknowledgment
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix H.
@@ Appendix C Details of the Subsets Division
As observed in Figure 9, subsets with high semantic shifts and low covariate shifts (or vice versa) contain few images, sometimes even none. This indicates that there is a certain correlation between semantic and covariate information in ImageNet-21K. Therefore, we have to sacrifice the results on some subsets (marking the results of subsets with too little data as "N/A"), in which way the levels of shift can cover a broader range of shift degrees (a greater shift degree in the highest shift level). Notably, the correlation between semantic and covariate information in ImageNet-21K could potentially lead to incorrect evaluations of OOD detection models that are sensitive to covariate shifts. By creating a more balanced dataset like Syn-IS, the bias caused by this correlation can be mitigated.
@@ F.5 AUPR Results on Syn-IS
Appendix G Further Analysis of Labeling Issues
Leveraging the classification model trained on ImageNet-1K, we analyze the prediction results for the images presented in Figure 2, along with the confidence scores derived using MSP [1], as illustrated in Figure 20. As observed, the model does indeed assign high confidence (i.e., low OOD scores) to the marginal OOD samples, highlighting the limitation of relying solely on semantic labels to distinguish OOD samples.
Figure 20: The prediction results and the MSP [1] confidence scores for the marginal OOD images.
We also conduct a quantitative analysis of the labeling issues. We compare the performance of the OOD detection models on the entire ImageNet-21K (with labeling issues) with its performance on subsets with larger semantic shifts (with fewer labeling issues), as shown in Table 8. The results show that models perform better on data with larger semantic shifts. (MSP with AUROC of 79.5 on data with larger semantic shifts and 68.0 on the entire ImageNet-21K.) This result is consistent with the findings in NINCO [14] (manually removes the objects with the labeling issue), where researchers noted that model performance improves when ID-objects are removed from the test set.
Table 8: Results of mean AUROC(%) on subsets with varying degrees of labeling issues
with labeling issues fewer labeling issues
MSP [1] 68.0 79.5
ODIN [2] 74.1 84.9
MDS [3] 52.3 58.9
GradNorm [4] 70.1 73.0
KNN [5] 65.4 78.7
DICE [6] 69.4 78.9
RankFeat [7] 57.3 59.0
ASH [8] 72.3 83.7
Appendix H Computational Resources
@@
Appendix I Definition of Semantic and Covariate Shift
Here, we give a formal definition of the semantic shift and the covariate shift mentioned in our work. We first define the semantic feature $s$ and the covariate feature $c$ for a sample, assuming that an image $x$ is determined by both $s$ and $c$, while the image label $y$ is influenced only by $s$:
$x$ $=f(s,c),\quad s\in S,c\in C\rightarrow x\in X,$ (7)
$y$ $=g(s),\quad s\in S\rightarrow y\in Y.$
We then provide the definitions for samples with semantic shift and covariate shift.
Semantic Shift:
$x^{ss}$ $=f(s^{\prime},c),\quad s^{\prime}\notin S,c\in C\rightarrow x^{ss}\notin X,$ (8)
$y^{ss}$ $=g(s^{\prime}),\quad s^{\prime}\notin S\rightarrow y^{ss}\notin Y.$
Covariate Shift:
$x^{cs}$ $=f(s,c^{\prime}),\quad s\in S,c^{\prime}\notin C\rightarrow x^{cs}\notin X,$ (9)
$y^{cs}$ $=g(s),\quad s\in S\rightarrow y^{cs}\in Y.$
In our benchmark, the semantic shift levels measure the distance between the test semantic $s^{\prime}$ and the ID semantic set $S$. The same applies to the covariate shift levels.
It’s worth noting that we observe ImageNet-21K includes labels such as beach and snow, which might serve as backgrounds (covariate contents) in other datasets, but appear as semantic contents in ImageNet-21K.
We treat all the labels in ImageNet-21K (such as backgrounds) as semantic contents in the proposed IS-OOD, as defined in Eq. 7, which might differ from some other benchmarks. In our experiments, the covariate contents only include styles, image augmentations, and lighting changes. However, future researchers can modify the collection of covariate texts according to their needs to study different elements as the covariate components.
Appendix J Future Directions
In terms of improving the benchmark, a promising direction for future research is to explore the use of more advanced vision-language models, such as BLIP or CoCa, to reduce the gap between text and image feature space, thereby enhancing the accuracy of feature decomposition in the LAID method. Additionally, separate investigations of particular covariate shifts are valuable, such as specifically studying the impact of changes in artistic styles on OOD detection model performance. We also strongly encourage subsequent researchers to use our benchmark for further evaluation of models in other fields, such as assessing whether the understanding of image semantic information by the current large-scale vision-language models, such GPT-4, is affected by image covariate shifts.
Appendix K Data Sheet
@@ Appendix I Definition of Semantic and Covariate Shift
K.4 Preprocessing/cleaning/labeling